(-5,-5), 6x-y=1,
Write an equation that passes through the given point and is parallel to the given line.

Answer by mananth(16946)

6 x + -1 y = 1
Find the slope of this line
make y the subject
-1 y = -6 x + 1
Divide by -1
y = 6 x -1
Compare this equation with y=mx+b
slope m = 6 0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be 6
m= 6 ,point ( -5 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = -30.00 + b
b= 35
m= 6
Plug value of the slope and b in y = mx +b
The required equation is y = 6 x + 35

m.ananth@hotmail.ca